OpenStudy (anonymous):

How would I solve this??? 4x+2=x+8

OpenStudy (anonymous):

First, group your like terms, 4x+2-2=x+8-2 = 4x=x+6 = 4x-x=x-x+6 = 3x=6 then divide to find X

OpenStudy (anonymous):

ty so the answer would be 4 right?? @Shadowkatt28

OpenStudy (usukidoll):

2...

OpenStudy (usukidoll):

all variables go to the left all numbers go to the right

OpenStudy (anonymous):

Explain??

OpenStudy (usukidoll):



OpenStudy (usukidoll):